OpenStudy (anonymous):

5x +y =6 4x +5y=2 which one is parallel, perpendicular or are neither

OpenStudy (anonymous):

re-write equations as: y=-5x+6 y==x*4/5 + 1/2 compare numbers in a front of x... neiher

OpenStudy (anonymous):

both equations are written in Standard Form: Ax + By = C the A for the first equation is 5 and the B is 1, so A/B = 5/1 the A for the 2nd equation is 4 and the B is 5, so A/B = 4/5 if A/B for both are the same, the lines are parallel. if the A/B for both multiply to make -1, they are perpendicular if neither is true, then the lines intersect and they are neither par or perp. check with your teacher, s/he might want you to solve for y like the above post mentions. But use my rule with the above post.